Calculate the strength in g/L of 3N HCl and N/2 H2SO4 solutions.

Grade:12th pass

1 Answers

Durvesh
6 years ago
Strength=Normality × Equivalent weightStrength (HCl)=3×36.5=109.5 g/lStrength (H2SO4)=1/2×98/2=98/4=24.5 g/l